OpenStudy (anonymous):

Use partial fractions to find the sum of the series 1/4n^2-1

OpenStudy (anonymous):

Use \[ \frac{1}{4 n^2-1} =\frac{1}{2 (2 n-1)}-\frac{1}{2 (2 n+1)} \] and use telescoping.

OpenStudy (anonymous):

Can you remind me of telescoping? Doesnt that mean that most terms will cancel each other out??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

You should find that the sum of the series is \[\frac 1 2 \]

OpenStudy (anonymous):

You have to show your work

OpenStudy (anonymous):

First factor the 1/2 out as \[ \frac 1 2 \left(\frac{1}{2 n-1}-\frac{1}{ 2 n+1}\right) \] and work with \[ a_n =\frac{1}{2 n-1}-\frac{1}{ 2 n+1} \] Notice that \[ a_1 = 1 - \frac 1 3\\ a_2 =\frac 13 -\frac 1 5\\ \] write \[ a_3=? -? \] then write \[a_{n-2}=?-?\\ a_{n-1}=?-?\\ a_{n}=?-?\\ \] Add all these to get \[s_n=\sum_{i=1}^n a_i \] and deduce that \[ s_n \] converges to 1. Do not forget the 1/2 we pulled out.

OpenStudy (anonymous):

I understand, thanks!!